leila1234567:

Solve the inequality and graph the solution. 2(y+7)≤6

Five:

brake it apart

leila1234567:

ok so 2y+7

jhonyy9:

so like a first step divide both sides by 2

leila1234567:

2 divide by 7 and 2 divided by 6?

jhonyy9:

no both sides by 2 so \[\frac{ 2(y+7) }{ 2 } <= \frac{ 6 }{ 2 }\]

leila1234567:

oh

jhonyy9:

simplifie and what will get ?

leila1234567:

y=4.5+3?

jhonyy9:

no in the left part simplifie 2 by 2 and what remain there ?

leila1234567:

1

leila1234567:

1y+7?

jhonyy9:

yes but this is just y+7 ok ? and in the other side ?

leila1234567:

y+7=3?

jhonyy9:

not equal y+7 <= 3

leila1234567:

oh ok then

jhonyy9:

and now how you isolate the y in the left part ? what you need to do ?

leila1234567:

so 3-7

leila1234567:

y<4?

jhonyy9:

yes subtract 7 from both sides and will get what ?

leila1234567:

4

jhonyy9:

so in this way in the left part remain y yes but in the right part 3-7 = ?

leila1234567:

-4?

jhonyy9:

yes -4 but between left part and right part not is equal sign

leila1234567:

y<-4 so the sign flips?

jhonyy9:

not is true bc. not is just <

leila1234567:

oh ok

jhonyy9:

so how will get y ?

leila1234567:

1 divided by -4?

jhonyy9:

Leila pls you wrote above that y<-4 but there are two signs < and = so how will be y ?

leila1234567:

so y < -4 but the inequlity sing?

jhonyy9:

Leila please check it again you missed from there a sign bc. y is smaller and equal or not ?

leila1234567:

ok

jhonyy9:

so how will be ?

leila1234567:

y>-4?

jhonyy9:

not really y is smaller and equal -4 so y <= -4 how will graph this ?

leila1234567:

so a closed point on -4 going to the <

jhonyy9:

yes closed point on -4 what mean that y can get the value of -4 and going to where ?

leila1234567:

to the left side

jhonyy9:

till the end of so what mean this mathematical ? in math words

leila1234567:

negative side

jhonyy9:

have you heard about infinity ?

leila1234567:

yes

jhonyy9:

so than how will be in the left part what infinity ?

leila1234567:

wait I'm confused

jhonyy9:

there are two infinity yes ? minus in the left part and plus on the right part in this your problem what will be ?

leila1234567:

-?

jhonyy9:

yes of corse courage pls

jhonyy9:

so what values can get y ?

leila1234567:

only negative past -4 or -4

jhonyy9:

y <= -4 so this mean that y can get values from -4 ,inclusive -4 till minus infinity ok ?

leila1234567:

ok

jhonyy9:

questions ?

leila1234567:

No and thank you

jhonyy9:

ok was my pleasure and anytime bye bye

leila1234567:

bye